PartnerObserving Log for 2007 Oct 16-17

Observers: D. Sand and M. Coleman
Telescope Operator: A. Ceranski

-- JohnHill - 17 Oct 2007

Sand and Coleman here.

0053 -- As sun is setting, there are angry looking clouds directly overhead. But I think they will pass soon.

New actuator has been put in and last minute tests are being performed. Have not taken biases because of this.

Windy outside. ~12 m/s

0204 -- Beginning 25 biases after everything has been worked out. Going outside to see if there are any stars..

0206 -- No stars.

0246 -- Went outside. Still all clouds..

0308 -- Another check. Still all clouds.

0342 -- Decided to open. Seems to be clearing up...windy.

0354 -- Noticing that bias frames look bad. A light is on in the stairwell in the dome.

0410 -- Light has finally been tracked down. Another bias frame shows that this was not the culprit.

0413 -- Talking to Hill. He says cycle power. Possibly several times. He has seen this problem before.

0421 -- Hill is the man. Cycling power works.

0422 -- Moving to Schwope's rbs1774_b_10

0429 -- Something strange is going on. rbs1774_b_10 is running, but we are not point at the right coordinates AT ALL. Bottom line -- Calling Hill

Our coordinates were not getting sent to the telescope, it times out, and LBC took an image -- as far as I can tell and I think that is what Hill is saying too.

0437 -- We think we got it solved. Moving to rbs1774_b_10, really.

0449 -- After a 16s and 32s Vfastextra with nothing in the images, we're giving up on this field. Outside is cloudy...

Moving to Clementini's M31_S2B_27 No luck.

0504 -- Continuing to take Vfastextra's while cooking and waiting for something to show up.

0551 -- Had to close due to high humidity

0856 -- Humidity at 88.2%. Sky is clear. We are opening the dome.

0901 -- Moving to M31_S2B_27. May be in vain. Humidity back to 90.5%.

0904 -- Vfastextra

0935 -- Finally got the focus figured out. Starting M31_S2B_27.

Well, we may be doomed. The mirror and ambient temperature are 5 degrees different. Our seeing from the guider output seems to be ~3-4 arcsec.

Looking at the first image, seeing is damn near 3 arcseconds. Going to a Ryan field.

0944 -- Going to Ryan. Oct_10_deg_epoch_1

0946 -- Vfastextra

0955 -- Starting Oct_10_deg_epoch_1

First image has nearly 3 arcsec seeing.

1009 -- Going to Ryan 20 deg field.

Vfastextra

1021 -- beginning Oct_20_deg_epoch_1

First image has ~2 arcsec imaging.

seeing is slowly improving...

second image has similar seeing.

1034 -- Moving to Oct_0_deg_epoch_1

Vfastextra

1040 -- Beginning Oct_0_deg_epoch_1

First image is 2.0" seeing.

Second image is 1.7" seeing.

I will party on with Ryan until they hit 1.6 AM.

1057 -- Vfastextra on Oct_10_deg_epoch_2

1103 -- Beginning Oct_10_deg_epoch_2

Ugh. LBC Control System GUI has frozen. Clicking on the Power Control tab and then returning to the OB Execution tab does not help. waiting for exposure to end, which is the next thing to try accoding to the troubleshooting page

[Note added later by JMH: Appears that LBC lost (misplaced?) the connection with TCS at 11:03:49. Log message is: watchdog on StatusThread started a new instance. (repeating every 20 sec) No corresponding messages on the TCS side until Aaron restarted CSQ at 11:14 UT.]

1133 -- Ok, had to call John Hill. Stopped and restarted lbc server with lbckill/lbcstart at 11:28. Just getting back online. The Ryan field is now at 1.8 AM.

1137 -- To be honest, don't know what to do. Seeing is 2", we think. Ryan fields are down. No great options. We are going to go to a Kochanek field and hope that the seeing is now 1.5 arcsec or something.

1141 -- Running Vfastextra version for M81

1150 -- Beginning KochanekM81BV. Seeing is 2+ arcsec according to guider. Bummer.

1153 -- After this I'll take some standards even though the night was definitely not photometric.

1204 -- Going to SA98_4p_UsUBVgr

1221 -- RPC Client/Host Error. Something wrong.

Trying another OB.

We have to reboot CMU according to trouble shooting page. It looks like our night is over!!

1237 -- Still restarting things...

1240 -- Aaron is slewing to FF spot, just in case I can get LBC back in time...

1241 SkyFlatG_5

1251 -- Running SkyFlatGrot180_5

Good night!
Topic revision: r5 - 21 Jul 2009, NormCushing
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback